# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ Abstract
Deep learning, a branch of artificial intelligence, is a data-driven method that uses multiple layers of interconnected units (neurons) to learn intricate patterns and representations directly from raw input data. Empowered by this learning capability, it has become a powerful tool for solving complex problems and is the core driver of many groundbreaking technologies and innovations. Building a deep learning model is challenging due to the algorithm’s complexity and the dynamic nature of real-world problems. Several studies have reviewed deep learning concepts and applications. However, the studies mostly focused on the types of deep learning models and convolutional neural network architectures, offering limited coverage of the state-of-the-art deep learning models and their applications in solving complex problems across different domains. Therefore, motivated by the limitations, this study aims to comprehensively review the state-of-the-art deep learning models in computer vision, natural language processing, time series analysis and pervasive computing. We highlight the key features of the models and their effectiveness in solving the problems within each domain. Furthermore, this study presents the fundamentals of deep learning, various deep learning model types and prominent convolutional neural network architectures. Finally, challenges and future directions in deep learning research are discussed to offer a broader perspective for future researchers.
@@
Deep Learning , Fundamentals , State-of-the-art Applications , Challenges
journal: Neurocomputing
@@ 1 Introduction
Deep learning has revolutionized many applications across a variety of industries and research. The application of deep learning can be found in healthcare [1], smart manufacturing [2], robotics [3], cybersecurity [4] etc., solving challenging and complex problems such as disease diagnosis, anomaly detection, object detection and malware attack detection. Deep learning is a subset of machine learning that focuses on learning from data using artificial neural networks with many layers, known as deep neural networks. An artificial neural network is a computational model that imitates the working principles of a human brain. The computational models are composed of an input layer which receives the input data, multiple processing layers that learn the representation of data and the output layer which produces the output of the model.
Prior to the reintroduction of deep learning (DL) into the research trend, pattern recognition tasks involved a transformation of the raw input data such as pixel values of an image into a feature vector that represents the internal representation of the data. The feature vector can be used by a machine learning model to detect or classify patterns in the data. This process requires feature engineering and considerable domain knowledge to design a suitable feature representation. With deep learning, this cumbersome process can be performed automatically whereby at each processing layer known as hidden layers, the internal representation of the input data is learned or extracted in a hierarchical manner. The first layer learns the presence of basic primitive features such as edges, dots, lines etc. The second layer learns patterns or motifs by recognizing the combinations of the edges, dots and lines, and the subsequent layers combine the motifs to produce more sophisticated features that correspond to the input data. This feature learning process takes place in the sequence of hidden layers until the prediction is finally produced.
@@
Despite the existing surveys on deep learning that offer valuable insights, the increasing amount of deep learning applications and the existing limitations in the current studies motivated us to explore this topic in depth. In general, to the best of our knowledge, no survey paper focuses on the emerging trends in state-of-the-art applications and the current challenges associated with deep learning. Furthermore, the surveys do not discuss the issues and how deep learning addresses them by highlighting the key features and components in the models. Also, most surveys either ignore or provide minimal coverage of the fundamentals of deep learning, which is crucial for understanding the state-of-the-art models. The main objective of this paper is to present the most important aspects of deep learning, making it accessible to a wide audience and facilitating researchers and practitioners in advancing and leveraging its capabilities to solve complex problems across diverse domains. Specifically, we present the fundamentals of deep learning and the various types of deep learning models, including popular deep learning architectures. Then, we discuss the progress of deep learning in state-of-the-art applications, highlighting the key features of the models and their problem-solving approaches. Finally, we discuss the challenges faced by deep learning and the future research directions.
@@
Reference Focus Concepts not covered
@@ 2.1 Layers
where $x_{j}$ is the input vector of size $d$, $w_{i,j}^{l}$ is the weight associated with $x_{j}$, connecting the input to neuron $i$ at layer $l$, and $g$ is the transformation function also known as activation function. It is worth noting that $w_{0}$ is called bias and $x_{0}=1$. A hidden layer wherein each neuron is connected to all neurons of the previous layer is known as fully-connected layer. The forward pass computation of a neuron $a_{i}^{l}$ in layer $l$ receiving a set of input from layer $l-1$ can be generalized as follows:
@@
Another crucial layer type is the convolutional layer, which is primarily used for processing data that has underlying structures, such as spatial patterns in image data or temporal patterns in time series data. Unlike fully-connected layer, each neuron in a convolutional layer is connected only to a subset of neurons in the previous layer as shown in Fig. 2. As shown in the figure, each neuron in the hidden layer is connected only to a local region, a subset of nine input neurons, and the weights are shared across the input data. The weight-sharing not only significantly reduces the number of parameters of the neural network, but also allows the network to learn the same features across different spatial locations in the input [11]. Let $\mathbf{x}$ be a two-dimensional input data with $H\times W\times C$ such as a grayscale image, where $H$ is the height, $W$ is the width and $C$ is the channel of the image data. The computation of a neuron in convolutional layer $l$ is defined as
@@
where $w_{m,n,d}$ is the weight connecting the input data $x_{i,j,d}$ within the window to the neuron in layer $l$ and $s$ is the stride or the step size of the window as it moves across the input data. This convolution operation where the filter is represented by $w_{m,n,d}$ slides over the input data $x_{i,j,d}$, producing a set of output values called feature map.
@@ 2.2 Attention Mechanisms
One of the important concepts in pattern recognition is the ability to attend and neglect certain parts of the input data based on their importance. This is because not all parts of the input hold equal importance for making the prediction. Certain features exhibit a stronger correlation with the output while others are less relevant. In convolutional layers, all extracted features are treated uniformly, without consideration of the varying degree of the importance of the different parts of the input data. This limitation is addressed by the introduction of attention mechanism, which can dynamically assign varying levels of significance (weights) to the different features. This flexibility enables the deep learning models to prioritize the more relevant aspects of the input data, enhancing its ability to capture the intricate dependencies for accurate prediction. Given an input data $x$, the process of attending to the important components of the input is given as
@@
For instance, the squeeze-and-excitation (SE) attention generates the attention through five consecutive operations [12]. First, the input is vectorized using global average pooling. Then the vector is passed to two fully-connected layers, where the first one with ReLU activation and the second one with sigmoid activation. SE attention was a pioneer in channel attention. The attention module assigns varying weights to the channels of the feature maps. SE attention suffers from computational cost and the use of global average which may cause information loss at the spatial level. Several efforts have been made to improve SE attention. Global Second-order Pooling (GSoP) attention performs $1\times 1$ convolution on the feature maps to reduce the number of channels, and then computes the pairwise channel correlation which is used to generate the weights [13]. Efficient Channel Attention (ECA) replaced the fully-connected layers with 1d convolution to reduce the number of parameters and the computational cost [14].
@@
Temporal attention is an attention module that focuses on specific time steps in a sequence of data such time series and video (sequence of images). In video processing such as recognizing human actions, temporal attention is used to focus on key frames at different point in time that contains crucial information for predicting the ongoing activity. Temporal adaptive module (TAM) is a temporal attention that can focus on short-term (local) information and global context information of the data [15]. The composite function consists of local branch for generating attention weights and global branch for generating channel-wise adaptive kernel. First, the input feature map is squeezed using global average pooling to reduce the computational cost. Subsequently, the local branch executes two 1D convolution operations, with the first convolution using ReLU, and sigmoid activation for the second to generate the local weights. The local weights are then multiplied with the featuremap. Meanwhile, the global branch is composed of two fully-connected layers, with the first layer using ReLU and second layer employing softmax function to generate the adaptive kernel (weights). Self-attention is a form of temporal attention, initially proposed for machine translation to enable the deep learning models to attend different words in a sequence relative to other words [16]. The attention module has become a fundamental building block in various natural language processing applications. To generate the attention weights, the input (word embeddings) is transformed by linear projection to compute query, key and value. Then, dot product between query and key is computed, and the resultant is normalized by the square root of the size of the key. Finally, the attention weights are obtained by applying the softmax function. Self-attention is the fundamental building block of the Transformer architecture, a key deep learning model in natural language processing.
@@
Spatial attention focuses on specific regions or spatial location of the input data, enabling the deep learning models to selectively emphasize and ignore certain features. In the context of computer vision, spatial attention is crucial in capturing the spatial relationships and context within an image for accurate prediction. Attention gate is a spatial attention that can identify and focus the salient regions and suppress feature responses of the insignificant ones. The composite function consists of ReLU activation followed by $1\times 1$ convolution to reduce channel dimension of the feature maps to a singular feature map. Finally, sigmoid is applied to the feature map to generate attention weights [17]. The self-attention in the standard Transformer is not effective in handling image data due to its inherent sequential processing nature and lacks the ability to capture spatial dependencies and local patterns. To address this limitation, the Vision Transformer (ViT) treats images as a sequence of non-overlapping patches. A similar computational pipeline is used to generate the attention weights, the sequence of patches is transformed by linear projection to compute the query, key and value [18]. The same operations are employed to generate the attention weights. Self-attention is computationally costly due to its quadratic complexity especially dealing with image data. To reduce the complexity, two learnable linear layers, independent of the input data are adopted as the key and value vectors [19].
@@ 2.3 Activation Functions
The role of the activation function is to transform the weighted sum into a more classifiable form. This is crucial to the learning behaviour of the deep learning model, generating non-linear relationships between the input and the output of the model. The activation function, combined with many hidden layers allow the neural network to approximate highly complex, non-linear functions. Many activation functions are available for use in neural networks, and some of the functions are shown in Fig. 4 - Fig 6. The figures show the plot of the three popular activation functions. The sigmoid is a classic example of activation function which is used in logistic regression. Sigmoid activation function maps the weighted sum to a value in the range of 0 and 1 which can be used for classification. Hyperbolic tangent is another popular choice of bounded activation function which produces an output between -1 and 1. Hyperbolic tangent has a stronger gradient, hence the neural network training often converges faster than sigmoid [20]. For many years, sigmoid and hyperbolic tangent are the commonly used activation functions. Nevertheless, the activation functions suffer from vanishing gradient problem, hindering the efficient training of deep neural networks with many layers [21].
It was shown that neural networks with unbounded activation functions have the universal approximation property and reduce the vanishing gradient problem. Over the past years, numerous unbounded activation functions were proposed for neural networks, with softplus [22] and rectified linear unit (ReLU) [23] activation function being the notable examples. These activation functions especially ReLU has pivotal role in improving the training and performance of deep learning models. ReLU has been a cornerstone of deep learning models due to its computational efficiency and effectiveness in addressing the vanishing gradient problem. Since then, several variants of ReLU were proposed including Leaky ReLU [24], sigmoid linear unit [25] and exponential linear unit [26], each offering unique advantages for building deep learning-based applications.
@@ 2.4 Parameter Learning and Loss Functions
The weights (parameters) of deep learning models are often optimized using an optimization algorithm called gradient descent, though other optimization algorithms may also be used. However, it has to be noted that gradient descent is a generic algorithm which can be used to solve a wide range of optimization problems. In general, gradient descent finds the optimal weights by iteratively updating the weights such that the weights will result in a minimum prediction error over all instances in the training set. The prediction error is quantified by a loss function. For classification problems, the commonly used loss function is the negative log-likelihood loss or cross entropy loss, while the square loss and absolute loss are used for regression problems [27]. The weights of layer $l$ is updated as
@@
The gradient can be computed across all training set instances, an approach known as batch gradient descent. However, this approach does not always guarantee convergence to the optimal solution, as it may get stuck in local minima or saddle points, and the same gradient is used for every weight update. An alternative approach is to perform the weight update on the basis of a single instance, but the approach results in a noisy gradient and becomes computationally intensive due to the frequent weight update. A more commonly used approach is to perform the weight update over a set of training instances, known as mini-batch gradient descent. This approach strikes a balance, providing a less noisy gradient and a more stable training process.

### v6

@@ Abstract
Deep learning, a branch of artificial intelligence, is a data-driven method that uses multiple layers of interconnected units or neurons to learn intricate patterns and representations directly from raw input data. Empowered by this learning capability, it has become a powerful tool for solving complex problems and is the core driver of many groundbreaking technologies and innovations. Building a deep learning model is challenging due to the algorithm’s complexity and the dynamic nature of real-world problems. Several studies have reviewed deep learning concepts and applications. However, the studies mostly focused on the types of deep learning models and convolutional neural network architectures, offering limited coverage of the state-of-the-art deep learning models and their applications in solving complex problems across different domains. Therefore, motivated by the limitations, this study aims to comprehensively review the state-of-the-art deep learning models in computer vision, natural language processing, time series analysis and pervasive computing, and robotics. We highlight the key features of the models and their effectiveness in solving the problems within each domain. Furthermore, this study presents the fundamentals of deep learning, various deep learning model types and prominent convolutional neural network architectures. Finally, challenges and future directions in deep learning research are discussed to offer a broader perspective for future researchers.
@@
Deep Learning , Fundamentals , Computer Vision , Natural Language Processing , Time Series Analysis , Robotics , Challenges
journal: Engineering Applications of Artificial Intelligence
@@ 1 Introduction
Deep learning has revolutionized many applications across a variety of industries and research. The application of deep learning can be found in healthcare [1], smart manufacturing [2], robotics [3] and cybersecurity [4], solving challenging and complex problems such as disease diagnosis, anomaly detection, object detection and malware attack detection. Deep learning is a subset of machine learning that focuses on learning from data using artificial neural networks with many layers, known as deep neural networks. An artificial neural network is a computational model that imitates the working principles of a human brain. The computational models are composed of an input layer which receives the input data, multiple processing layers that learn the representation of data and the output layer which produces the output of the model.
Prior to the reintroduction of deep learning (DL) into the research trend, pattern recognition tasks involved a transformation of the raw input data such as pixel values of an image into a feature vector that represents the internal representation of the data. The feature vector can be used by a machine learning model to detect or classify patterns in the data. This process requires feature engineering and considerable domain knowledge to design a suitable feature representation. With deep learning, this cumbersome process can be performed automatically whereby at each processing layer known as hidden layers, the internal representation of the input data is learned or extracted in a hierarchical manner. The first layer learns the presence of basic primitive features such as edges, dots, and lines. The second layer learns patterns or motifs by recognizing the combinations of the edges, dots and lines, and the subsequent layers combine the motifs to produce more sophisticated features that correspond to the input data. This feature learning process takes place in the sequence of hidden layers until the prediction is finally produced.
Deep learning has seen numerous breakthroughs in various industries, transforming how complex problems are solved and how businesses operate. For instance, protein folding is a complex process, influenced by the intricate sequence of amino acids. Using traditional methods to determine how a protein folds would take an immense amount of computational power and time due to its complex 3D structure and vast number of possible configurations. AlphaFold [5] has made significant progress in this area, using deep learning to predict protein structures with remarkable accuracy, helping researchers to comprehend the complete structure of a key protein associated with diseases like malaria and Parkinson’s disease. In natural language processing, systems like OpenAI GPT, Google Gemini and IBM Watsonx have revolutionized chatbots and virtual assistants, enabling them to understand and respond to human language with remarkable accuracy and contextual awareness. These systems have overcome the challenge of processing vast amounts of unstructured text, allowing them to engage in more natural conversations and handle a wide range of topics. Businesses can use these systems to automate customer service, provide real-time support and improve user experience.
@@
Despite the existing surveys on deep learning that offer valuable insights, the increasing amount of deep learning applications and the existing limitations in the current studies motivated us to explore this topic in depth. In general, to the best of our knowledge, no survey paper focuses on the emerging trends in state-of-the-art applications and the current challenges associated with deep learning. Furthermore, the surveys do not discuss the issues and how deep learning addresses them by highlighting the key features and components in the models. Furthermore, most surveys either ignore or provide minimal coverage of the fundamentals of deep learning, which is crucial for understanding the state-of-the-art models. The main objective of this paper is to present the most important aspects of deep learning, making it accessible to a wide audience and facilitating researchers and practitioners in advancing and leveraging its capabilities to solve complex problems across diverse domains. Specifically, we present the fundamentals of deep learning and the various types of deep learning models, including popular deep learning architectures. Then, we discuss the progress of deep learning in state-of-the-art applications, highlighting the key features of the models and their problem-solving approaches. Finally, we discuss the challenges faced by deep learning and the future research directions.
To this end, we conducted automatic search strategies using “deep learning” keyword with keywords related to deep learning applications such as “image classification”, “neural machine translation”, “text generation”, “human activity recognition” and “robotics”. Then, a manual search was carried out by scanning the references produced by the automatic search, selecting relevant works and discarding the irrelevant ones. This survey paper collected primary studies from journals, conference proceedings, and books in English only. We use scientific databases such as IEEE Explore, ScienceDirect, SpringerLink, ACM Digital Library, Scopus and ArXiv. The ArXiv repository is used because it hosts numerous manuscripts that are highly relevant to the topics discussed in this survey. These preprints often contain novel findings and methodologies that may not be available in other repositories at the time of writing. Nevertheless, we ensure that the selected preprints are of high quality, authored by researchers from reputable institutions.
@@
Ref. Focus Concepts not covered
@@ 2.1 Layers
where $x_{j}$ is the input vector of size $d$, $w_{i,j}^{l}$ is the weight associated with $x_{j}$, connecting the input to neuron $i$ at layer $l$, and $g$ is the transformation function also known as activation function. It is worth noting that $w_{0}$ is called bias and $x_{0}=1$. A hidden layer wherein each neuron is connected to all neurons of the previous layer is known as the fully connected layer. The forward pass computation of a neuron $a_{i}^{l}$ in layer $l$ receiving a set of input from layer $l-1$ can be generalized as follows:
@@
Another crucial layer type is the convolutional layer, which is primarily used for processing data that has underlying structures, such as spatial patterns in image data or temporal patterns in time series data. Unlike a fully connected layer, each neuron in a convolutional layer is connected only to a subset of neurons in the previous layer, as shown in Fig. 2. As shown in the figure, each neuron in the hidden layer is connected only to a local region, a subset of nine input neurons, and the weights are shared across the input data. The weight-sharing not only significantly reduces the number of parameters of the neural network, but also allows the network to learn the same features across different spatial locations in the input [12]. Let $\mathbf{x}$ be a two-dimensional input data with $H\times W\times C$ such as a grayscale image, where $H$ is the height, $W$ is the width and $C$ is the channel of the image data. The computation of a neuron in the convolutional layer $l$ is defined as
@@
where $w_{m,n,d}$ is the weight connecting the input data $x_{i,j,d}$ within the window to the neuron in layer $l$ and $s$ is the stride or the step size of the window as it moves across the input data. This convolution operation where the filter is represented by $w_{m,n,d}$ slides over the input data $x_{i,j,d}$, producing a set of output values called feature map. For a more hands-on understanding of this convolutional operation, readers are encouraged to explore CNN Explainer [13], an interactive tool that demonstrates how convolutional layers process input data and generate feature maps.
@@ 2.2 Attention Mechanisms
One of the important concepts in pattern recognition is the ability to attend to and neglect certain parts of the input data based on their relevance. This is because not all parts of the input hold equal importance for making predictions. Certain features exhibit a stronger correlation with the output while others are less relevant. To provide a simple analogy, consider a set of sensors to measure room temperature. The sensors are deployed across different corners of the room. Each sensor measures temperature in its own area, and these local measurements are used to estimate the room temperature. However, not all sensors are equally important due to various factors. For instance, a sensor near an air-conditioning or a heater might give inaccurate readings due to the external temperature fluctuation caused by these systems. A sensor near the center of the room, away from the heating or cooling sources might provide more reliable readings. Therefore, when estimating the room temperature, we give more attention to the sensors that are less affected by the localized factors. In convolutional layers, all extracted features are treated uniformly, without consideration of the varying degree of the importance of the different parts of the input data. This limitation is addressed by the introduction of attention mechanism, which can dynamically assign varying levels of significance (weights) to the different features. This flexibility enables the deep learning models to prioritize the more relevant aspects of the input data, enhancing its ability to capture the intricate dependencies for accurate prediction. Given an input data $x$, the process of attending to the important components of the input is given as
@@
For instance, the squeeze-and-excitation (SE) attention generates the attention through five consecutive operations [14]. First, the input is vectorized using global average pooling. Then the vector is passed to two fully connected layers, where the first one with ReLU activation and the second one with sigmoid activation. SE attention was a pioneer in channel attention. The attention module assigns varying weights to the channels of the feature maps. SE attention suffers from computational cost and the use of global average, which may cause information loss at the spatial level. Several efforts have been made to improve SE attention. Global Second-order Pooling (GSoP) attention performs $1\times 1$ convolution on the feature maps to reduce the number of channels, and then computes the pairwise channel correlation, which is used to generate the weights [15]. Efficient Channel Attention (ECA) replaced the fully connected layers with 1D convolution to reduce the number of parameters and the computational cost [16].
@@
Temporal attention is an attention module that focuses on specific time steps in a sequence of data such as time series and video (sequence of images). In video processing such as recognizing human actions, temporal attention is used to focus on key frames at different point in time that contain crucial information for predicting the ongoing activity. Temporal adaptive module (TAM) is a temporal attention that can focus on short-term (local) information and global context information of the data [17]. The composite function consists of a local branch for generating attention weights and a global branch for generating a channel-wise adaptive kernel. First, the input feature map is squeezed using global average pooling to reduce the computational cost. Subsequently, the local branch executes two 1D convolution operations, with the first convolution using ReLU, and sigmoid activation for the second to generate the local weights. The local weights are then multiplied with the feature map. Meanwhile, the global branch is composed of two fully connected layers, with the first layer using ReLU and the second layer employing softmax function to generate the adaptive kernel (weights). Self-attention is a form of temporal attention, initially proposed for machine translation to enable the deep learning models to attend different words in a sequence relative to other words [18]. The attention module has become a fundamental building block in various natural language processing applications. To generate the attention weights, the input (word embeddings) is transformed by linear projection to compute query, key, and value. Then, the dot product between query and key is computed, and the resultant is normalized by the square root of the size of the key. Finally, the attention weights are obtained by applying the softmax function. Self-attention is the fundamental building block of the Transformer architecture, a key deep learning model in natural language processing. For a more hand-on understanding of self-attention and Transformer, readers are encouraged to explore Transformer Explainer [19], an interactive tool that demonstrates how self-attention mechanisms work in Transformer by visualizing the attention scores and how different inputs interact with each other.
@@
Spatial attention focuses on specific regions or spatial location of the input data, enabling the deep learning models to selectively emphasize and ignore certain features. In the context of computer vision, spatial attention is crucial in capturing the spatial relationships and context within an image for accurate prediction. Attention gate is a spatial attention that can identify and focus the salient regions and suppress feature responses of the insignificant ones. The composite function consists of ReLU activation followed by $1\times 1$ convolution to reduce channel dimension of the feature maps to a singular feature map. Finally, sigmoid is applied to the feature map to generate attention weights [20]. The self-attention in the standard Transformer is not effective in handling image data due to its inherent sequential processing nature and lacks the ability to capture spatial dependencies and local patterns. To address this limitation, the Vision Transformer (ViT) treats images as a sequence of non-overlapping patches. A similar computational pipeline is used to generate the attention weights, the sequence of patches is transformed by linear projection to compute the query, key, and value [21]. The same operations are employed to generate the attention weights. Self-attention is computationally costly due to its quadratic complexity, especially when dealing with image data. To reduce the complexity, two learnable linear layers, independent of the input data, are adopted as the key and value vectors [22].
@@ 2.3 Activation Functions
The role of the activation function is to transform the weighted sum into a more classifiable form. This is crucial to the learning behavior of the deep learning model, generating non-linear relationships between the input and the output of the model. The activation function, combined with many hidden layers, allows the neural network to approximate highly complex, non-linear functions. Many activation functions are available for use in neural networks, and some of the functions are shown in Fig. 4—Fig 6. The figures show the plots of the three popular activation functions. The sigmoid is a classic example of an activation function used in logistic regression. It maps the weighted sum to a value in the range of 0 to 1 which can be used for classification. The hyperbolic tangent (tanh) is another popular choice of bounded activation function, which produces an output between -1 and 1. Since it has a stronger gradient, neural network training often converges faster than with the sigmoid function [23]. For many years, sigmoid and hyperbolic tangent functions were the commonly used activation functions. However, both suffer from the vanishing gradient problem, which hinders the efficient training of deep neural networks with many layers [24].
It was shown that neural networks with unbounded activation functions have the universal approximation property and reduce the problem of vanishing gradients. In recent years, numerous unbounded activation functions have been proposed for neural networks, with the softplus [25] and the rectified linear unit (ReLU) [26] activation function being notable examples. These activation functions, especially ReLU has a pivotal role in improving the training and performance of deep learning models. ReLU has been a cornerstone of deep learning models due to its computational efficiency and effectiveness in addressing the vanishing gradient problem. Since then, several variants of ReLU have been proposed including Leaky ReLU [27], sigmoid linear unit [28] and exponential linear unit [29], each offering unique advantages for building deep learning-based applications.
@@ 2.4 Parameter Learning and Loss Functions
The weights (parameters) of deep learning models are often optimized using an optimization algorithm called gradient descent, though other optimization algorithms may also be used. However, it has to be noted that gradient descent is a generic algorithm which can be used to solve a wide range of optimization problems. In general, gradient descent finds the optimal weights by iteratively updating the weights such that the weights will result in a minimum prediction error over all instances in the training set. This process is synonymous with a person at the top of a hill who wants to climb down to the ground. Just like the person may choose a path that leads to the lowest point by taking small steps based on the steepness of the slope, gradient descent makes small adjustments to the weights, moving them in the direction that reduces the error the most. The prediction error is quantified by a loss function. For classification problems, the commonly used loss function is the negative log-likelihood loss or cross entropy loss, while the square loss and absolute loss are used for regression problems [30]. The weights of layer $l$ are updated as
@@
The gradient can be computed across all training set instances, an approach known as batch gradient descent. However, this approach does not always guarantee convergence to the optimal solution, as it may get stuck in local minima or saddle points, and the same gradient is used for every weight update. An alternative approach is to perform the weight update on the basis of a single instance, but the approach results in a noisy gradient and becomes computationally intensive due to the frequent weight update. A more commonly used approach is to perform the weight update over a set of training instances, known as mini-batch gradient descent. This approach strikes a balance, providing a less noisy gradient and a more stable training process. For a more hands-on understanding of neural network training, readers are encouraged to explore TF playground [31] and Initializing Neural Networks [32], interactive tools demonstrating how neural networks are trained and the impact of hyperparameters.